OpenStudy (anonymous):

Hi, can you help me find the inflection point(s) of this function f(x)=(x^3)/((x^2)-4) I found -2, 0 and 2 as my answer but it's wrong. Please help Thank you :)

OpenStudy (solomonzelman):

inflection points are where f''(x) =0.

OpenStudy (solomonzelman):

I don't see any way to simplify the f(x) for the first derivative, so I would just go ahead and use a quotient rule. Or, re-write the x^2-4 thas on the bottom, as (x^2-4)^(-1) and then differentiate it with product rule.

OpenStudy (solomonzelman):

what was your first derivative?

OpenStudy (anonymous):

Actually I found my second derivative with online calculator

OpenStudy (anonymous):

...

OpenStudy (solomonzelman):

why calculator, why not by hand?

OpenStudy (anonymous):

Because it's too difficult for me to find the second derivative.

OpenStudy (solomonzelman):

It is not that difficult.

OpenStudy (solomonzelman):

you can use the product rule for second derivative.

OpenStudy (solomonzelman):

I mean re-writing the bottom as (x^2-4)^(-2)

OpenStudy (solomonzelman):

The point is that you know HOW to find the derivative, so do you know how to find the first one without a calculator?

OpenStudy (anonymous):

I think I do, but I always get a wrong answer. So really no I don't

OpenStudy (solomonzelman):

Maybe because you are getting the rules incorrectly:) \(\LARGE\color{black}{ \frac{d}{dx}\left(\begin{matrix} \frac{F(x)}{G(x)} \\ \end{matrix}\right) =\frac{g(x)~\color{red}{f \prime (x)}~-\color{red}{g \prime (x)}~f (x)}{ \left[ ~g(x)~ \right] ^2 } }\)

OpenStudy (solomonzelman):

lets do it with your function, now,

OpenStudy (solomonzelman):

tell me first 2 derivatives: 1) x^2-4 2) x^3 okay?

OpenStudy (anonymous):

1) 2x = g'(x) 2) 3x^2 = f'(x)

OpenStudy (solomonzelman):

good.

OpenStudy (solomonzelman):

I got disconnected again

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{d}{dx}\left(\begin{matrix} \frac{\color{blue}{f(x)}}{\color{red}{g(x)}} \\ \end{matrix}\right) =\frac{\color{red}{g(x)}~\color{blue}{f \prime (x)}~-\color{red}{g \prime (x)}~\color{blue}{f (x)}}{ \left[ ~\color{red}{g(x)}~ \right] ^2 } }\) \(\LARGE\color{black}{ \frac{d}{dx}\left(\begin{matrix} \frac{\color{blue}{x^3}}{\color{red}{x^2-4}} \\ \end{matrix}\right) =\frac{\color{red}{(x^2-4)}~\color{blue}{3x^2}~-\color{red}{2x}~\color{blue}{(x^3)}}{ \left[ ~\color{red}{(x^2-4)}~ \right] ^2 } }\)

OpenStudy (solomonzelman):

Can you simplify this?

OpenStudy (solomonzelman):

no need simplifying the bottom, just do the top part.

OpenStudy (solomonzelman):

you can just draw it, (as long as I can read it).

OpenStudy (solomonzelman):

or type the simplified result for the top that you get, here.

OpenStudy (anonymous):

\[(3x^4-12x^2-2x^4)/(x^2-4)^2\]

OpenStudy (solomonzelman):

yes, and what cancels on the top? subtract like terms, and you get?

OpenStudy (anonymous):

x^2

OpenStudy (solomonzelman):

where is this coming from?

OpenStudy (solomonzelman):

jjst x^2 on the top??

OpenStudy (solomonzelman):

*just

OpenStudy (anonymous):

No wait

OpenStudy (solomonzelman):

k

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

Because 3x^4-2x^4

OpenStudy (anonymous):

x^4-12x^2

OpenStudy (solomonzelman):

yes, exactly, so, \(\LARGE\color{black}{ \frac{d}{dx}\left(\begin{matrix} \frac{\color{blue}{x^3}}{\color{red}{x^2-4}} \\ \end{matrix}\right) =\frac{x^4-12x^2 }{ \left[ ~\color{red}{(x^2-4)}~ \right] ^2 } }\)

OpenStudy (solomonzelman):

now we need to find the second derivative.

OpenStudy (solomonzelman):

tell me first (again same thing) the derivative off: 1) x^4 -12x^2 2) (x^2-4)^2 .

OpenStudy (anonymous):

should we be using the quotient rule and chain rule?

OpenStudy (solomonzelman):

we will have to use chain rule in either case, but we have a choice betwen a quotient rule, or product rule with negative exponenets.

OpenStudy (anonymous):

1) 4x^3-24x 2) 2x*2(x^2-4)

OpenStudy (solomonzelman):

yes. that is right.

OpenStudy (solomonzelman):

We will use a product rule okay?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{d}{dx} \left[\color{white}{\frac{a}{b}}(x^4-12x^2)~(x^2-4)^{-2}~ \right] }\) this is what we will need to find.

OpenStudy (solomonzelman):